# Normalize Filters! Classical Wisdom for Deep Vision

### v3

@@ 4.3 Analysis and ablation studies
We investigate the relationship between a filter’s level of “unnormalization" and misclassifications. First, we sort the filters in the first layer of an ImageNet pretrained R34 based on the absolute value of their positive weight ratio ($|r(w)|$). We expect this metric to be 0 for normalized differencing filters (See Fig. 14left). Then, focusing on the first layer, where gain and bias effects are largest, we find the Top 100 images on $D_{C}$ that maximize the response for each filter and calculate the number of misclassified ones by the network.
@@
$|r(w)|$ Filters ranked by $|r(w)|$
Figure 13: Unnormalized filters misclassify more often. We investigate how filter weight normalization affects error rates in a pretrained R34. Left: We show the histogram of $|r(w)|$ from all filter weights in the network. We want this ratio to be 0 if the filters are differencing, or 1 if they are averaging filters. The peak in 0 from our R34N indicates all our filters are differencing and normalized, while vanilla ResNets possess many unnormalized. Right: We sort filters by “unnormalization” level and find that images triggering strong responses from “less normalized” filters (i.e., higher $|r(w)|$) are more often misclassified. Our convolutions avoid this by design.
@@ A.7 Further experimental validation
Lastly, we further assess generalization to natural variation by binning the ImageNet-1k test set into 9 contrast levels (based on pixel standard deviation). As shown in Table A9, R34N maintains more stable accuracy across the contrast spectrum compared to the baseline, which drops sharply at extreme values. This confirms the model’s improved robustness to natural realistic atmosphere conditions.
Table A9: Accuracy per contrast level. We bin the ImageNet-1k test set into 9 contrast levels. R34N maintains more stable accuracy across the contrast spectrum compared to the baseline, which drops sharply at extreme values.

### v4

@@ 4.1 Results on artificially corrupted benchmarks
Robustness on original ImageNet images across low and high contrast.
To evaluate robustness to natural variation, we partition the uncorrupted ImageNet-1k test set into nine contrast bins defined by pixel-level standard deviation. As shown in Table 2, R34N exhibits notably more consistent accuracy across contrast levels, while the baseline model suffers substantial degradation at the extremes. These results highlight the improved robustness of R34N to natural, realistic atmospheric conditions.
Table 2: Accuracy per contrast level. We bin the ImageNet-1k test set into 9 contrast levels (based on pixel standard deviation). R34N maintains more stable accuracy across the contrast spectrum compared to the baseline, which drops sharply at the extreme contrast values. Top row: example images from the uncorrupted test set for the lowest-contrast bin (left), the middle bin (center), and the highest-contrast bin (right). Bottom: accuracy table showing performance for each contrast bin.
@@ 4.3 Analysis and ablation studies
We investigate the relationship between a filter’s level of “unnormalization" and misclassifications. First, we calculate the positive weight ratio ($|r(w)|$) of the filters in the first layer of an ImageNet pretrained R34. We expect this metric to be 0 for normalized differencing filters (See Fig. 14left). Then, focusing on the first layer, where gain and bias effects are largest, we find the Top 100 images on $D_{C}$ that maximize the response for each filter and calculate the number of misclassified ones by the network.
@@
$|r(w)|$ $|r(w)|$
Figure 13: Unnormalized filters misclassify more often. We investigate how filter weight normalization affects error rates in a pretrained R34. Left: We show the histogram of $|r(w)|$ from all filter weights in the network. We want this ratio to be 0 if the filters are differencing, or 1 if they are averaging filters. The peak in 0 from our R34N indicates all our filters are differencing and normalized, while vanilla ResNets possess many unnormalized. Right: We find that images triggering strong responses from “less normalized” filters (i.e., higher $|r(w)|$) are more often misclassified. Our convolutions avoid this by design.